# Effect of Off-diagonal NSI Parameters on Entanglement Measurements in Neutrino Oscillations

### v1

@@ III.1 Entanglement of Formation
In quantum mechanics, a pure state is said to be entangled if it cannot be written as a product of states of its subsystems. That is, the total state $\gamma$ cannot be expressed as $\gamma_{A}\otimes\gamma_{B}$, then the state is entangled. A convenient way to measure this entanglement is through the entropy of entanglement, which is defined as the von Neumann entropy of either of the reduced density matrices: Bennett et al. (1996)
@@
where $S(\rho)=-\text{Tr}(\rho\log_{2}\rho)$ is the von Neumann entropy. The reduced density matrix $\rho_{A}$ is obtained by tracing out subsystem $B$, $\rho_{A}=\text{Tr}_{B}\left(\ket{\gamma}\bra{\gamma}\right)$, and vice versa for $\rho_{B}$.
When dealing with mixed states, entanglement can still be quantified using the EOF. For a mixed state, EOF, $M$ is defined as the minimum average entanglement over all possible pure-state decompositions that realize $M$ Bennett et al. (1996):
$E(M)=\min\sum_{i}p_{i}E\left(\gamma_{i}\right),$
where $M=\sum_{i}p_{i}\ket{\gamma_{i}}\bra{\gamma_{i}}$ | and the minimization is over all possible sets of pure states |$\ket{\gamma_{i}}$ and corresponding probabilities $p_{i}$ that reproduce the mixed state $M$.
@@
where $\rho_{A}$ and $\rho_{B}$ are the reduced density matrices of the two subsystems. This definition can also be extended to tripartite systems. For a pure three-partite state $\rho_{ABC}(t)$, the EOF is given by Guo and Zhang (2020):
@@
where $P_{\alpha\beta}$ are the oscillation probabilities from $\nu_{\alpha}$ to $\nu_{\beta}$ transition.
@@ III.2 Concurrence
For oscillating neutrinos, the concurrence for vacuum can be expressed in terms of oscillation and survival probabilities as Li et al. (2021)
@@
where $P_{S}$ is the sum of the square of probabilities, and defined as $(P_{\mu e}^{2}+P_{\mu\mu}^{2}+P_{\mu\tau}^{2})$.
@@ III.3 Negativity
In terms of survival and oscillation probabilities of neutrino oscillation, negativity is given as Li et al. (2021) :
@@
IV Experimental setup and simulation details
In this analysis, we consider the upcoming long-baseline neutrino experiment DUNE (Deep Underground Neutrino Experiment). DUNE is an ambitious forthcoming project featuring a baseline of 1300 km, extending from the Fermi National Accelerator Laboratory (FNAL) to the Sanford Underground Research Facility (SURF). For simulation purposes, we utilize the official configuration files based on the technical design report Abi et al. (2021). The experimental setup includes four liquid argon time-projection chamber (LArTPC) detectors, each with a mass of 10 kilotons, and operates with a 1.2 MW proton beam. The files represent an exposure of 624 kt-MW-years which corresponds to 6.5 years of run each in neutrino (FHC) and antineutrino (RHC) modes, using a 40 kt fiducial mass liquid argon time-projection chamber (LArTPC) far detector. For systematic errors, we use the numerical values from the ref. Abi et al. (2021).
To perform the DUNE simulations, we employ the General Long-Baseline Experiment Simulator (GLoBES) framework Huber et al. (2005, 2007). The probability engine within GLoBES has been modified to incorporate the effects of non-standard interactions (NSI) in the neutrino oscillation probability calculations.
@@ IV Experimental setup and simulation details
To estimate the sensitivities, we use the Poisson log-likelihood method, defined as
$\chi^{2}=2\sum_{i=1}^{n}\left[\mathcal{N}_{i}^{\rm test}-\mathcal{N}_{i}^{\rm true}-\mathcal{N}_{i}^{\rm true}\rm{log}\left(\frac{\mathcal{N}_{i}^{\rm test}}{\mathcal{N}_{i}^{\rm true}}\right)\right],$ (14)
where, $\mathcal{N}_{i}^{\rm true}$ and $\mathcal{N}_{i}^{\rm test}$ are the event numbers in the true and test spectra, respectively, and the index $i$ runs over the energy bins. The true values of the oscillation parameters used in the analysis are taken from Ref. Esteban et al. (2020). A marginalization over the relevant oscillation parameters is performed during the $\chi^{2}$ analysis. In this work, we restrict our results to the normal mass ordering scenario, i.e., $\Delta m_{31}^{2}>0.$
V Results
In this section, we will illustrate the main results of our simulation. In the first subsection, we show how the appearance and disappearance probabilities are modified in the presence of off-diagonal NSI parameters, incorporating one parameter at a time. Then, in the following subsection, we discuss the effect of off-diagonal NSI parameters in the entanglement measures: EOF, Concurrence, and Negativity.
@@ V.1 Analysis in probability level
In this subsection, we discuss how non-standard interaction (NSI) parameters influence the appearance and disappearance probabilities within the framework of the DUNE experiment. At the source, a beam of muon neutrinos ($\nu_{\mu}$) is generated and travels through matter over a baseline of 1300 km to reach the far detector. Therefore, our primary focus is on the oscillation ($P_{\mu e}$) and survival ($P_{\mu\mu}$) probabilities. The presence of NSI modifies the standard probability expressions. In this analysis, we concentrate specifically on the off-diagonal NSI parameters: $\epsilon_{e\mu}$, $\epsilon_{e\tau}$, and $\epsilon_{\mu\tau}$. Since these are off-diagonal elements, they can possess both real and imaginary components, each associated with a corresponding phase: $\phi_{e\mu}$, $\phi_{e\tau}$, and $\phi_{\mu\tau}$. The appearance probability $P_{\mu e}$, modified by the presence of off-diagonal NSI parameters, can be expressed as follows  [^1] In Ref. Liao et al. (2017), the expression for $P_{\mu e}$ also includes contributions from diagonal NSI terms (such as $\epsilon_{ee}$). However, since our analysis focuses exclusively on the off-diagonal parameters, Eq. V.1 is formulated solely in terms of these off-diagonal contributions, keeping $\epsilon_{ee}=0$. Liao et al. (2017):
@@
$\quad+4\hat{A}\epsilon_{e\mu}\Big{\{}xf[s_{23}^{2}f\cos(\phi_{e\mu}+\delta_{CP})+c_{23}^{2}g\cos(\Delta+\delta_{CP}+\phi_{e\mu})]$
@@
$\quad+4\hat{A}\epsilon_{e\tau}s_{23}c_{23}\Big{\{}xf[f\cos(\phi_{e\tau}+\delta_{CP})-g\cos(\Delta+\delta_{CP}+\phi_{e\tau})]$
@@
$\quad+8\hat{A}^{2}fgs_{23}c_{23}\Big{\{}c_{23}\cos\Delta[s_{23}(\epsilon_{e\mu}^{2}-\epsilon_{e\tau}^{2})+2c_{23}\epsilon_{e\mu}\epsilon_{e\tau}\cos(\phi_{e\mu}-\phi_{e\tau})]$
$\quad-\epsilon_{e\mu}\epsilon_{e\tau}\cos(\Delta-\phi_{e\mu}+\phi_{e\tau})\Big{\}}+\mathcal{O}(s_{13}\epsilon,s_{13}\epsilon^{2},\epsilon^{3}),$ (15)
@@
for normal ordering. For the inverted mass ordering, the corresponding substitutions are: $\Delta\rightarrow-\Delta,y\rightarrow-y$ and $\hat{A}\rightarrow-\hat{A}$. In the case of antineutrinos, Eq. V.1 is modified by the following transformations: $\hat{A}\rightarrow-\hat{A},\delta_{CP}\rightarrow-\delta_{CP}$ and $\phi_{\alpha\beta}\rightarrow-\phi_{\alpha\beta}$. Here $c_{ij}=\cos\theta_{ij}$, $s_{ij}=\sin\theta_{ij}$ with $i,j=1,2,3$ and $A=2\sqrt{2}G_{F}N_{e}E$, is the standard matter potential. $G_{F}$ is the Fermi constant and $N_{e}$ is the electron number density. From Eq. V.1, it is evident that $P_{\mu e}$ largely depends on the off-diagonal NSI parameters $\epsilon_{e\mu}$, $\epsilon_{e\tau}$ and their associated phases, while it remains nearly independent of $\epsilon_{\mu\tau}$ (however, at subleading order, a dependence on $\epsilon_{\mu\tau}$ does appear).
The disappearance probability in presence of off-diagonal NSI parameters can be written as  [^2] Similar to V.1, in the Ref. Liao et al. (2017), the expression for $P_{\mu\mu}$ also includes contributions from diagonal NSI terms (such as $\epsilon_{\mu\mu},\epsilon_{\tau\tau}$). However, Eq. V.1 is formulated solely in terms of off-diagonal contributions, keeping $\epsilon_{\mu\mu}=\epsilon_{\tau\tau}=0$. Liao et al. (2017):
@@
$-2\hat{A}\epsilon_{\mu\tau}\cos\phi_{\mu\tau}\Big{[}\sin^{3}2\theta_{23}\Delta\sin 2\Delta+2\sin 2\theta_{23}\cos^{2}2\theta_{23}\sin^{2}\Delta\Big{]}$
$-2\hat{A}^{2}\sin^{2}2\theta_{23}\epsilon_{\mu\tau}^{2}\Big{[}2\sin^{2}2\theta_{23}\cos^{2}\phi_{\mu\tau}\Delta^{2}\cos 2\Delta+\sin^{2}\phi_{\mu\tau}\Delta\sin 2\Delta\Big{]}.$ (17)
Figure 1: Upper (lower) row shows the appearance probability as a function of neutrino energy for DUNE experiment with $\delta_{CP}=212^{\circ}(0^{\circ})$. Left (middle) [right] panel represent the curves with one off-diagonal parameter one at a time, $\epsilon_{e\mu}$ ($\epsilon_{e\tau}$) [$\epsilon_{\mu\tau}$] and their associated phases. Color codes are given in the legend.
@@
The left-most panel of the upper row in Fig. 1 displays the effect of the NSI parameter $\epsilon_{e\mu}$ on the appearance probability when we take $\delta_{CP}=212^{\circ}$. From the plot, it is observed that the green curve exhibits a higher amplitude than the red curve. In contrast, the blue curve shows a reduced amplitude compared to the red one. To understand this behavior, we refer to the modified expression for $P_{\mu e}$ in the presence of NSI, considering only $\epsilon_{e\mu}$ and $\phi_{e\mu}$ (i.e., setting $\epsilon_{e\tau}=0$). The corresponding appearance probability can be expanded as:
$P(\nu_{\mu}\to\nu_{e})|_{\epsilon_{e\tau,\phi_{e\tau}=0}}$ $=SO+D1+E1+F1+G1+J1+K1+L1,$ (18)
@@
$SO=x^{2}f^{2}+2xyfg\cos(\Delta+\delta_{CP})+y^{2}g^{2},~{}~{}~{}~{}~{}D1=4\hat{A}\epsilon_{e\mu}xf^{2}s_{23}^{2}\cos(\phi_{e\mu}+\delta_{CP}),$
$E1=4\hat{A}\epsilon_{e\mu}xfgc_{23}^{2}\cos(\phi_{e\mu}+\Delta+\delta_{CP}),~{}~{}~{}~{}~{}~{}~{}~{}F1=4\hat{A}\epsilon_{e\mu}yg^{2}c_{23}^{2}\cos(\phi_{e\mu}),$
$G1=4\hat{A}\epsilon_{e\mu}ygfs_{23}^{2}\cos(\Delta-\phi_{e\mu}),~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}J1=4\hat{A}^{2}g^{2}\epsilon_{e\mu}^{2}c_{23}^{2},$
$K1=4\hat{A}^{2}f^{2}\epsilon_{e\mu}^{2}s_{23}^{4},~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}L1=8\hat{A}^{2}fg\epsilon_{e\mu}^{2}\sin\theta_{23}\cos\theta_{23}(\cos\theta_{23}\cos\Delta\sin\theta_{23}).$ (19)
Figure 2: Upper (lower) row shows the effect of each term containing $\epsilon_{e\mu}$ and $\phi_{e\mu}$ in appearance probability as a function of neutrino energy with $\delta_{CP}=212^{\circ}~{}(0^{\circ})$. Left (right) panel shows the results with $\phi_{e\mu}=0~{}(1.38\pi$). Color codes are given in the legend. Each term is thoroughly explained within the text.
The left and right panels of upper row of Fig. 2 depict $P_{\mu e}$ as a function of neutrino energy for $\epsilon_{e\mu}=0.15$ with $\phi_{e\mu}=0$ and $1.38\pi$, respectively. In both cases, we show the individual contributions from each term in Eq. 18. It is evident that D1 and E1 terms dominate the overall modification to the appearance probability. When $\phi_{e\mu}=0$, the D1 term contributes negatively, while the E1 term gives a positive enhancement in amplitude up to around 4 GeV, followed by a reduction beyond that. Although these contributions act in opposite directions, their combined effect increases the overall amplitude without shifting the oscillation peak, since their maximum and minimum occur at the same energy.
For $\phi_{e\mu}=1.38\pi$, the behavior changes notably. The E1 term now contributes negatively across the entire energy range, and the D1 term becomes nearly negligible. This change in behavior of E1 term arises due to the cosine dependence in E1: $\cos(\phi_{e\mu}+\Delta+\delta_{CP})$. At $\delta_{CP}=212^{\circ}$ and $\phi_{e\mu}=0$, this term is positive around the oscillation peak. However, for $\phi_{e\mu}=1.38\pi$, it becomes negative over the full energy spectrum. A similar reasoning applies to the D1 term, which contains $\cos(\phi_{e\mu}+\delta_{CP})$ and thus has minimal contribution at $1.38\pi$. Additionally, in the upper row of the leftmost panel of Fig. 1, the green curve exceeds the red curve only up to about 4 GeV, after which the red curve overtakes. This behavior aligns with the left panel of the upper row of Fig. 2, where the E1 term becomes negative beyond 4 GeV for $\phi_{e\mu}=0$.
@@
Figure 4: The disappearance probability as a function of neutrino energy for the DUNE experiment with $\delta_{CP}=212^{\circ}$. Left (middle) [right] panel represent the curves with one off-diagonal parameter one at a time, $\epsilon_{e\mu}$ ($\epsilon_{e\tau}$) [$\epsilon_{\mu\tau}$] and their associated phases. Color codes are given in the legend.
Figure 5: The effect of each term containing $\epsilon_{\mu\tau}$ and $\phi_{\mu\tau}$ in disappearance probability as a function of neutrino energy with $\delta_{CP}=212^{\circ}$. Left (right) panel shows the results with $\phi_{\mu\tau}=0~{}(0.6\pi$ rad). Color codes are given in the legend. For more about each term, see the text.
Unlike the green curve in Fig. 1, the blue curve consistently shows a lower amplitude than the red curve across the full energy range up to 8 GeV. This suppression is directly attributable to the negative E1 contribution shown in the right panel of upper row of Fig. 2, which dominates when $\phi_{e\mu}=1.38\pi$.
@@
To investigate the combined effect of NSI and the CP-violating phase $\delta_{CP}$ on the appearance probability, we consider a CP-conserving scenario with $\delta_{CP}=0^{\circ}$, as shown in the lower row in Fig. 1. The color scheme in this panel follows the same convention as the upper row: the red curve corresponds to standard oscillation, the green curve represents the case with only $\epsilon_{e\mu}$, and the blue curve includes both $\epsilon_{e\mu}$ and the phase $\phi_{e\mu}$. From the leftmost plot of the lower row of Fig. 1, it is evident that the red curve lies below both the green and blue curves across the entire energy range. Between the green and blue curves, the green curve shows a higher amplitude than the blue one.
This behavior is further clarified in the lower panel of Fig. 2, which decomposes the individual contributions to $P_{\mu e}$ under the same parameter settings as the upper panel but with $\delta_{CP}=0^{\circ}$. Similar to the upper panels, the D1 and E1 terms are found to contribute most significantly to the overall modification of the appearance probability.
The overall amplitude for the configuration with $\epsilon_{e\mu}=0.15$ and $\phi_{e\mu}=0$ is larger than that of the setup with $\epsilon_{e\mu}=0.15$ and $\phi_{e\mu}=1.38\pi$; this trend is clearly visible in the lower panel of Fig. 2.

### v2

@@ III Entanglement measures
Entanglement is one of the most fundamental features of quantum mechanics, reflecting the existence of non-classical correlations between different parts of a composite system. A state is said to be entangled if it cannot be expressed as a simple product of the states of its individual subsystems, meaning that the description of each part is inseparable from the whole. Quantifying entanglement is essential not only for understanding fundamental aspects of quantum theory but also for exploring its role as a resource in quantum communication, computation, and information processing. To capture its different facets, several entanglement measures have been introduced, each offering a unique perspective on the strength and nature of quantum correlations. In this work, we focus on three widely employed measures: Entanglement of Formation, Concurrence, and Negativity and derive these measures in the context of neutrino oscillation probabilities.
@@ III.1 Entanglement of Formation
In quantum mechanics, a pure state is said to be entangled if it cannot be written as a product of states of its subsystems. That is, when the total state $\gamma$ cannot be expressed as $\gamma_{A}\otimes\gamma_{B}$, then the state is entangled. A convenient way to measure this entanglement is through the entropy of entanglement, which is defined as the von Neumann entropy of either of the reduced density matrices: Bennett et al. (1996)
@@
[…]
When dealing with mixed states, entanglement can still be quantified using the EOF. For a mixed state, EOF is defined as the minimum average entanglement over all possible pure-state decompositions that realize $M$ Bennett et al. (1996):
$EOF(M)=\min\sum_{i}p_{i}E\left(\gamma_{i}\right),$
[…]
@@
where $\rho_{A}$ and $\rho_{B}$ are the reduced density matrices of the two subsystems. This definition can also be extended to tripartite systems. For a pure tripartite system, the density matrix is denoted as $\rho_{ABC}(t)$ and the EOF is given by Guo and Zhang (2020):
@@
where $P_{\alpha\beta}$ are the oscillation probabilities from $\nu_{\alpha}$ to $\nu_{\beta}$ transition. The detailed derivation of the above expression of the Entanglement of Formation (EOF) in terms of oscillation probabilities is provided in Appendix A.
@@ III.2 Concurrence
For oscillating neutrinos, the Concurrence for $\mu$ as initial flavor, $C^{\mu}$ can be expressed in terms of oscillation and survival probabilities as Li et al. (2021)
@@
where $P_{S}$ is the sum of the square of probabilities, and defined as $(P_{\mu e}^{2}+P_{\mu\mu}^{2}+P_{\mu\tau}^{2})$. The derivation of the Concurrence expression in terms of neutrino oscillation probabilities can be found in Appendix A.
@@ III.3 Negativity
In terms of survival and oscillation probabilities of neutrino oscillation, Negativity for $\mu$ as initial flavor, $N^{\mu}$ is given as Li et al. (2021) :
@@
The explicit derivation of the Negativity formula, expressed through oscillation and survival probabilities, is presented in Appendix A.
IV Experimental Description
In this analysis, we consider the upcoming long-baseline neutrino experiment Deep Underground Neutrino Experiment (DUNE). DUNE is an ambitious forthcoming project featuring a baseline of 1300 km, extending from the Fermi National Accelerator Laboratory (FNAL) to the Sanford Underground Research Facility (SURF). For simulation purposes, we utilize the official configuration files based on the technical design report Abi and others (2021). The experimental setup includes four liquid argon time-projection chamber (LArTPC) detectors, each with a mass of 10 kilotons, and operates with a 1.2 MW proton beam. The files represent an exposure of 624 kt-MW-years which corresponds to 6.5 years of run each in neutrino (FHC) and antineutrino (RHC) modes, using a 40 kt fiducial mass liquid argon time-projection chamber (LArTPC) far detector. In this work, we restrict our results to the normal mass ordering scenario, i.e., $\Delta m_{31}^{2}>0.$
@@ IV Experimental Description
V Results
In this section, we will illustrate the main results of our simulation. In the first subsection, we show how the appearance and disappearance probabilities are modified in the presence of off-diagonal NSI parameters, incorporating one parameter at a time. We then extend this study to the event level in the following subsection. Then, in the next subsection, we discuss the effect of off-diagonal NSI parameters in the entanglement measures: EOF, Concurrence, and Negativity. The final subsection shows the variation of entanglement measures with neutrino energy and CP-violating phase, $\delta_{CP}$.
@@ V.1 Analysis in probability level
In this subsection, we discuss how non-standard interaction (NSI) parameters influence the appearance and disappearance probabilities within the framework of the DUNE experiment. At the source, a beam of muon neutrinos ($\nu_{\mu}$) is generated and travels through matter over a baseline of 1300 km to reach the far detector. Therefore, our primary focus is on the oscillation ($P_{\mu e}$) and survival ($P_{\mu\mu}$) probabilities. The presence of NSI modifies the standard probability expressions. In this analysis, we concentrate specifically on the off-diagonal NSI parameters: $\epsilon_{e\mu}=\left|\epsilon_{e\mu}\right|e^{-i\phi_{e\mu}}$, $\epsilon_{e\tau}=\left|\epsilon_{e\tau}\right|e^{-i\phi_{e\tau}}$, and $\epsilon_{\mu\tau}=\left|\epsilon_{\mu\tau}\right|e^{-i\phi_{\mu\tau}}.$ Since these are off-diagonal elements, they can possess both real and imaginary components, each associated with a corresponding phase: $\phi_{e\mu}$, $\phi_{e\tau}$, and $\phi_{\mu\tau}$. The appearance probability $P_{\mu e}$, modified by the presence of off-diagonal NSI parameters, can be expressed as follows  [^1] In Ref. Liao et al. (2017), the expression for $P_{\mu e}$ also includes contributions from diagonal NSI terms (such as $\epsilon_{ee}$). However, since our analysis focuses exclusively on the off-diagonal parameters, Eq. V.1 is formulated solely in terms of these off-diagonal contributions, keeping $\epsilon_{ee}=0$. Liao et al. (2017, 2016b); Kopp et al. (2008); Kikuchi et al. (2009):
@@
$\quad+4\hat{A}\left|\epsilon_{e\mu}\right|\Big\{xf[s_{23}^{2}f\cos(\phi_{e\mu}+\delta_{CP})+c_{23}^{2}g\cos(\Delta+\delta_{CP}+\phi_{e\mu})]$
@@
$\quad+4\hat{A}\left|\epsilon_{e\tau}\right|s_{23}c_{23}\Big\{xf[f\cos(\phi_{e\tau}+\delta_{CP})-g\cos(\Delta+\delta_{CP}+\phi_{e\tau})]$
@@
$\quad+8\hat{A}^{2}fgs_{23}c_{23}\Big\{c_{23}\cos\Delta[s_{23}(\left|\epsilon_{e\mu}\right|^{2}-\left|\epsilon_{e\tau}\right|^{2})+2c_{23}\left|\epsilon_{e\mu}\right|\left|\epsilon_{e\tau}\right|\cos(\phi_{e\mu}-\phi_{e\tau})]$
$\quad-\left|\epsilon_{e\mu}\right|\left|\epsilon_{e\tau}\right|\cos(\Delta-\phi_{e\mu}+\phi_{e\tau})\Big\}+\mathcal{O}(s_{13}\epsilon,s_{13}\epsilon^{2},\epsilon^{3}),$ (21)
@@
for normal ordering. For the inverted mass ordering, the corresponding substitutions are: $\Delta\rightarrow-\Delta,y\rightarrow-y$ and $\hat{A}\rightarrow-\hat{A}$. In the case of antineutrinos, Eq. V.1 is modified by the following transformations: $\hat{A}\rightarrow-\hat{A},\delta_{CP}\rightarrow-\delta_{CP}$ and $\phi_{\alpha\beta}\rightarrow-\phi_{\alpha\beta}$. Here $c_{ij}=\cos\theta_{ij}$, $s_{ij}=\sin\theta_{ij}$ with $i,j=1,2,3$. From Eq. V.1, it is evident that $P_{\mu e}$ largely depends on the off-diagonal NSI parameters $\left|\epsilon_{e\mu}\right|$, $\left|\epsilon_{e\tau}\right|$ and their associated phases, while it remains nearly independent of $\epsilon_{\mu\tau}$ (however, at subleading order, a dependence on $\epsilon_{\mu\tau}$ does appear).
The disappearance probability in presence of off-diagonal NSI parameters can be written as  [^2] Similar to V.1, in the Ref. Liao et al. (2017), the expression for $P_{\mu\mu}$ also includes contributions from diagonal NSI terms (such as $\epsilon_{\mu\mu},\epsilon_{\tau\tau}$). However, Eq. V.1 is formulated solely in terms of off-diagonal contributions, keeping $\epsilon_{\mu\mu}=\epsilon_{\tau\tau}=0$. Liao et al. (2017, 2016b); Kopp et al. (2008); Kikuchi et al. (2009):
@@
$-2\hat{A}\left|\epsilon_{\mu\tau}\right|\cos\phi_{\mu\tau}\Big[\sin^{3}2\theta_{23}\Delta\sin 2\Delta+2\sin 2\theta_{23}\cos^{2}2\theta_{23}\sin^{2}\Delta\Big]$
$-2\hat{A}^{2}\sin^{2}2\theta_{23}\left|\epsilon_{\mu\tau}\right|^{2}\Big[2\sin^{2}2\theta_{23}\cos^{2}\phi_{\mu\tau}\Delta^{2}\cos 2\Delta+\sin^{2}\phi_{\mu\tau}\Delta\sin 2\Delta\Big].$ (23)
@@
In this analysis, we focus solely on the off-diagonal NSI parameters $\left|\epsilon_{\alpha\beta}\right|$ ($\alpha\neq\beta$), incorporating their associated complex phases $\phi_{\alpha\beta}$. To simplify the discussion, we consider one NSI parameter at a time. This allows us to examine the distinct effects of each off-diagonal NSI term on the oscillation probabilities. We explore three cases for each off diagonal NSI parameters $\epsilon_{\alpha\beta}$: (i) the standard oscillation scenario (SO) (ii) NSI with a vanishing complex phase $(\phi_{\alpha\beta}=0)$, and (iii) NSI with a non-zero complex phase $(\phi_{\alpha\beta}\neq 0)$. In each panel of Fig. 1, the red curve represents the standard oscillation (SO) scenario, while the green curve corresponds to the case with $\left|\epsilon_{\alpha\beta}\right|$ included. The blue curve illustrates the effect when both $\left|\epsilon_{\alpha\beta}\right|$ and its associated phase $\phi_{\alpha\beta}$ are taken into account. The shaded region corresponds to the DUNE $\nu_{\mu}$ neutrino flux in arbitrary units.
The left-most panel of the upper row in Fig. 1 displays the effect of the NSI parameter $\epsilon_{e\mu}$ on the appearance probability when we take $\delta_{CP}=212^{\circ}$. From the plot, it is observed that the green curve exhibits a higher amplitude than the red curve. In contrast, the blue curve shows a reduced amplitude compared to the red one. This variation in amplitudes can be understood by examining the modified expression for $P_{\mu e}$ in the presence of NSI. Considering only $\left|\epsilon_{e\mu}\right|$ and $\phi_{e\mu}$ (with $\left|\epsilon_{e\tau}\right|,$ and $\phi_{e\tau}=0$), the appearance probability can be expanded, and the individual contributions are explicitly displayed in Fig. 2 for a clearer understanding of the results in Fig. 1. Similarly, to further interpret the features of Fig. 1, we analyze the case with only $\left|\epsilon_{e\tau}\right|$ and $\phi_{e\tau}$ (with $\left|\epsilon_{e\mu}\right|,$ and $\phi_{e\mu}=0$), which is presented in Fig. 3.
Figure 1: Upper (lower) row shows the appearance probability as a function of neutrino energy for DUNE experiment with $\delta_{CP}=212^{\circ}(0^{\circ})$. Left (middle) [right] panel represent the curves with one off-diagonal parameter at a time, $\left|\epsilon_{e\mu}\right|$ ($\left|\epsilon_{e\tau}\right|$) $\left|\epsilon_{\mu\tau}\right|$] and their associated phases. Color codes are given in the legend.
To understand this behavior of Fig. 1, we refer to the modified expression for $P_{\mu e}$ in the presence of NSI, considering only $\left|\epsilon_{e\mu}\right|$ and $\phi_{e\mu}$ (i.e., setting $\left|\epsilon_{e\tau}\right|,\phi_{e\tau}=0$). The corresponding appearance probability can be expanded as:
$P(\nu_{\mu}\to\nu_{e})|_{\left|\epsilon_{e\tau}\right|},\phi_{e\tau}=0$ $=SO+D1+E1+F1+G1+J1+K1+L1,$ (24)
@@
$SO=x^{2}f^{2}+2xyfg\cos(\Delta+\delta_{CP})+y^{2}g^{2},~~~~~D1=4\hat{A}\left|\epsilon_{e\mu}\right|xf^{2}s_{23}^{2}\cos(\phi_{e\mu}+\delta_{CP}),$
$E1=4\hat{A}\left|\epsilon_{e\mu}\right|xfgc_{23}^{2}\cos(\phi_{e\mu}+\Delta+\delta_{CP}),~~~~~~~~F1=4\hat{A}\left|\epsilon_{e\mu}\right|yg^{2}c_{23}^{2}\cos(\phi_{e\mu}),$
$G1=4\hat{A}\left|\epsilon_{e\mu}\right|ygfs_{23}^{2}\cos(\Delta-\phi_{e\mu}),~~~~~~~~~~~~~~~~~J1=4\hat{A}^{2}g^{2}\left|\epsilon_{e\mu}\right|^{2}c_{23}^{2},$
$K1=4\hat{A}^{2}f^{2}\left|\epsilon_{e\mu}\right|^{2}s_{23}^{4},~~~~~~~L1=8\hat{A}^{2}fg\left|\epsilon_{e\mu}\right|^{2}\sin\theta_{23}\cos\theta_{23}(\cos\theta_{23}\cos\Delta\sin\theta_{23}).$ (25)
Figure 2: Upper (lower) row shows the effect of each term containing $\left|\epsilon_{e\mu}\right|$ and $\phi_{e\mu}$ in appearance probability as a function of neutrino energy with $\delta_{CP}=212^{\circ}~(0^{\circ})$. Left (right) panel shows the results with $\phi_{e\mu}=0~(1.38\pi$). Color codes are given in the legend. Each term is thoroughly explained within the text.
The upper row of Fig. 2 shows $P_{\mu e}$ as a function of neutrino energy for $\left|\epsilon_{e\mu}\right|$ = 0.15 with $\phi_{e\mu}=0$ and $1.38\pi$. Individual contributions from Eq. 24 reveal that D1 and E1 dominate. For $\phi_{e\mu}=0$, D1 is negative while E1 enhances the amplitude up to around 4 GeV, then decreases. Despite opposing signs, their combined effect raises the overall amplitude without shifting the oscillation peak, as their extrema occur at the same energy.
For $\phi_{e\mu}=1.38\pi$, the behavior changes notably. The E1 term now contributes negatively across the entire energy range, and the D1 term becomes nearly negligible. This arises from the cosine dependence in E1, $\cos(\phi_{e\mu}+\Delta+\delta_{CP})$, which is positive near the oscillation peak for $\phi_{e\mu}=0$ and $\delta_{CP}=212^{\circ}$ but becomes entirely negative at $\phi_{e\mu}=1.38\pi$. Similarly, the D1 term, proportional to $\cos(\phi_{e\mu}+\delta_{CP})$, contributes minimally at $1.38\pi$. Consistently, in Fig. 1 (upper-left panel), the green curve exceeds the red curve only up to $\sim$4 GeV, after which the red dominates, matching the trend in Fig. 2 where the E1 term turns negative beyond 4 GeV for $\phi_{e\mu}=0$. Unlike the green curve, the blue curve consistently shows a lower amplitude than the red curve for SO across the full energy range up to 8 GeV. This suppression is directly attributable to the negative E1 contribution shown in the right panel of the upper row of Fig. 2, which dominates when $\phi_{e\mu}=1.38\pi$.
@@
To investigate the combined effect of NSI and the CP-violating phase $\delta_{CP}$ on the appearance probability, we consider a CP-conserving scenario with $\delta_{CP}=0^{\circ}$, as shown in the lower row in Fig. 1. The color scheme in this panel follows the same convention as the upper row: the red curve corresponds to standard oscillation, the green curve represents the case with only $\left|\epsilon_{e\mu}\right|$, and the blue curve includes both $\left|\epsilon_{e\mu}\right|$ and the phase $\phi_{e\mu}$. From the leftmost plot of the lower row of Fig. 1, it is evident that the red curve lies below both the green and blue curves across the entire energy range. Between the green and blue curves, the green curve shows a higher amplitude than the blue one. This behavior is further clarified in the lower panel of Fig. 2, which decomposes the individual contributions to $P_{\mu e}$ under the same parameter settings as the upper panel but with $\delta_{CP}=0^{\circ}$. The D1 and E1 terms are found to contribute most significantly to the overall modification of the appearance probability. The overall amplitude for the configuration with $\left|\epsilon_{e\mu}\right|=0.15$ and $\phi_{e\mu}=0$ is larger than that of the setup with $\left|\epsilon_{e\mu}\right|=0.15$ and $\phi_{e\mu}=1.38\pi$; this trend is clearly visible in the lower panel of Fig. 2.